40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 5 07:19:16 2012, kiwamu, Update, LSC, DRMI locking DRMI_2012Apr4_edit.png
    Reply  Sat Apr 7 01:56:05 2012, kiwamu, Update, LSC, OSA signal in DRMI condition OSA.png
       Reply  Sat Apr 7 06:58:34 2012, kiwamu, Update, LSC, DRMI lock : lost good alignment 
Message ID: 6508     Entry time: Sat Apr 7 06:58:34 2012     In reply to: 6506
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: DRMI lock : lost good alignment 

Somehow I lost the good alignment, where the lock can be frequently acquired and hence I didn't go further ahead.

I will try locking the DRMI during the weekend again. My goal is to take time series when the DRMI is being locked and sensing matrix.

Quote from #6506

Currently SRM is slightly misaligned such that the MICH optical gain at AS55Q doesn't increase so much with the presence of SRM.

With this condition I was able to acquire the lock more frequently than how it used to be on the Wednesday.

The next step is to gradually align SRM, to optimize the controls and to repeat this process several times until SRM is fully aligned.

 

ELOG V3.1.3-